Timing Center - Maser Log

as of Sunday, 17-Dec-2017 17:38:22 AST

Go to Most Recent Entry


NameEmailDateComment
ANS anolla@naic.edu 12/19/2000 The HM was installed back as Household reference since 12/18/2000 AST 13:21.It had problems with the Vacion , WG fixed it. GPS 3 was been used as household reference while HM was out of service & the Rub 1 supplied the 5 Mhz for the Buffers.
ANS anolla@naic.edu 1/23/2001 The HM is out of service for maint. by Don Rhine , he is performing the Vacion replacement & adjustments on it.
ANS anolla@naic.edu 1/25/2001 HM back as household ref since 1/25/2001. Don Rhine replaced both Vacion pumps on unit.
ANS anolla@naic.edu 3/13/2001 W Genter replaced the back up batteries in the HM. Were 4 of them ,the replacement are Power Sonic model PS-12750 NBH 12 VDC 75 AH.
LW lwray@naic.edu 3/16/2001 The HM wheels are in the storage van. Also spare parts for the HM are in the cryo shop in a shelf especially for them.
LW lwray@naic.edu 3/19/2001 On 3/16/2001 WG & LW adjusted the T Time clock , now is +/- .140 usec.
LW lwray@naic.edu 4/20/2001 WG is working with the Stanford clock software. He found that the comp DB 9 connector was wrongly wired.
LW lwray@naic.edu 5/4/2001 HM charging system was not working properly, so batteries were weak (17 vdc instead 24 vdc) .Due to this clocks came out of sync when try to replace the UPS on HM. Clocks were reseted.
LW lwray@naic.edu 5/7/2001 Rub 1 & 2 were sync. Rub 1 is 9.72 us & Rub 2 is 8.07 us.
LW lwray@naic.edu 8/17/2001 Air Conditioner unit quit working since midnight , so auxiliary ac unit at window turned on by telescope operator. Room temp was at 87 f , doors opened to let cooled air circulate.
LW lwray@naic.edu 9/5/2001 HM failed due to the 18 VDC power supply , it was replaced with the Rub 1 standard since 9/4/2001. HM back into service
LW lwray@naic.edu 9/17/2001 Room temp at 80f due to failure on main AC unit. Window unit turned on.
LW lwray@naic.edu 2/26/2002 Room was without air conditioner from 1:00 AM to 8:00 AM , nobody turned on the auxiliary ac unit at the window until 8:00 AM.
ansanolla@naic.eduTuesday, August 27, 2002 13:51:03Temp chart paper recorder pens filled with ink.
ansanolla@naic.eduWednesday, November 13, 2002 08:55:12DATUM clock replacing the Rb 1 , is 20 nsec off from HM.
LSWlwray@naic.eduMonday, December 2, 2002 17:27:13100 MHz reference was out of lock, noticed on Thursday night (SETI observations). Crystal oscillator has been re-tuned. No idea yet how long this has been out of lock.
ansanolla@naic edu.Thursday, February 6, 2003 08:46:14NIST recvr was out during the last 5 days. Its 1PPS sync signal from HM was out for 718 msec. It happen since 1/30/2003 due to a power glitch.
ansanolla@naic edu.Tuesday, March 4, 2003 14:23:43The problem with the NIST was that the maser clock A was not sync with clock B. Have to double check it after reseting the HM.
ansanolla@naic.eduTuesday, March 4, 2003 14:26:42The HM battery charger was not working due to an open 10 amp fuse in the battery charger ckt. It was replaced and it is charging . This condition has more than six months with it.EC found the problem.
ansanolla@naic edu.Wednesday, March 5, 2003 12:52:49The HM was tested for battery auxiliary power.It's 110 vac was removed and was left with the battery back up , no problems with it.It worked perfectlly. The batteries were checked again before these test .baterry 1 was 12.9 v , battery 2 was 13.4 v ,battery 3 was 12.1 v and battery 4 was 12.6 v.
ansanolla@naic edu.Tuesday, March 11, 2003 12:33:38Batteries for the power back up been ordered by EC , They are Power Sonic model PS-12750 NHB 12VDC 75 AH.
ansanolla@naic.eduWednesday, March 19, 2003 14:51:22After a series of consecutive power dips check the clocks an WALLLAAAA!!!!!!! All of them on time.Thanks Eddie!!!!!!!
ansanolla@naic.eduFriday, March 28, 2003 09:01:14There it was a power dip that lasted around 1 minute.HM and clocks were ok except GPS 3 & Stanford clock that it's tide to it.
ansanolla@naic.eduTuesday, April 1, 2003 16:13:08Commercial power was out since AST 08:34 until AST 15:45.All clocks are OK except the GPS 3 that is out since power dip on 3/28/2003, Calixto is working on it.
ansanolla@naic.eduTuesday, April 8, 2003 12:44:11GPS 3 back into service.While checking power supply and power for the rf section of the GPS it came back functional.Why???? I don't know , but it is working.
ansanolla@naic.eduTuesday, April 22, 2003 13:18:44Ed Castro adjusted the cavity , it was set from 30676 to 31581 , it's a 905 bit increase. Was a Bill Diner request.
ansanolla@naic.eduTuesday, May 20, 2003 14:21:27Calixto was checking the phase jitter of the D. Backer GPS ,and to the Datum GPS.
ansanolla@naic.eduTuesday, May 27, 2003 14:53:06The 28 vdc main power supply was bad , it had a high ripple on it (.5 v ). Was replaced with the spare one , but was wired outside the HM due tha was not identical as the one on the HM.
ansanolla@naic.eduThursday, May 29, 2003 10:19:51The external main 28 vdc power supply was triped off forunknow reason. So the HM was running on its auxiliary battery supply.We do not want to remove the original power supply until next maintenance time in order to avoid a further problem with the HM.
ansanolla@naic.eduMonday, June 16, 2003 14:21:53The main + 28 power supply was fixed by IC , it had it's electolitic capacitors bad, all were replaced.
ansanolla@naic.eduMonday, June 16, 2003 14:25:28The Datum clock was removed in order to send it to factory for modification , a 5 mhz rub osc is going to be installed to have a 5 mhz out available.
ansanolla@naic.eduThursday, June 19, 2003 09:01:01The computer table was swapped in order to open an space for the new Datum HM that will arrive this Friday.
ansanolla@naic.eduMonday, June 23, 2003 13:15:13The new Datum HM arrived.Calixto is working on it , it came without batteries for the back up power.
ansanolla@naic.eduTuesday, July 1, 2003 10:10:39Panasonic batteries LC-X1242P installed in the new Datum HM.They are in series to have +24 vdc as an auxiliary power back up.
ansanolla@naic.eduThursday, July 3, 2003 08:47:49Calixto is performing some test with the new Datum HM. He is compearing the old HM 1PPS out with the new Datum HM.The T Time clock been used to produce the Datum 1 PPS , we locked it to the Datum using the 10 mhz out of it to lock the T Time clock ref oscillator.
ansanolla@naic.eduMonday, July 7, 2003 09:10:40The T Time clock was set to have the Datum 10 mhz ref on it , in oder words we are compearing the Old HM with the Datum HM thru the T Time clock 1 PPS.
ansanolla@naic.eduFriday, July 11, 2003 15:49:27The Datum clock is back after its modification ( add a 5 mhz output ) Calixto installed it.
ansanolla@naic.eduMonday, July 14, 2003 14:58:43The HM back up power supply batteries were replaced. Were Power Sonic PS-12750 12 vdc 75 AH. Now are YUASA DM55-12.
ansanolla@naic.eduMonday, August 11, 2003 15:32:54There is a new table and computer at the timing room.
ansanolla@naic.eduWednesday, August 13, 2003 08:50:22The Stanford clock was removed , then a TRAK clock model 9000A was installed instead. Calixto is working in its settings.
ansanolla@naic.eduWednesday, October 1, 2003 16:04:34The DATUM Maser was connected to the NIST receiver ,it will be like that for a month for testing purposes.
ansanolla@naic.eduWednesday, October 8, 2003 14:25:26Program not running.Was reset.There were power dips.
ansanolla@naic.eduMonday, December 1, 2003 16:26:37The NIST receiver was set up again with the old hydro Maser ( 1PPS & 5 Mhz). It was done at AST 14:15.
ansanolla@naic.eduWednesday, January 14, 2004 15:24:50The main 28 vdc power supply quit working so we put the spare one , but it didn't last.Power consumption of these P.S. is 5 amp and the power consumption of the HM ckt is above 8 amps. So we put a lab one that is up to 10 amps.CR sent for a spares P.S. Also the HM was reset by Ed Castro on 1/14/2003 at AST 14:00. All clocks are on time.
ansanolla@naic.eduTuesday, March 2, 2004 08:52:18Calixto tried to replace the power supply from the lab to the existing repaired one , but didn't work. Today (3/2/2004) the room temp is 4 higher than normal.
ansanolla@naic.eduTuesday, March 2, 2004 16:25:06Calixto tried the fix Lambda P.S. , but didn't work . HM load was tripping it off. So HM with lab P.S. for a while.
ansanolla@naic.eduWednesday, March 3, 2004 10:19:45 The HM old P.S. was installed after been fix , but it didn't worked .It was not holding the load , it is supposed that it's load is 8 amps.In the need of a new one P.S.
ansanolla@naic.eduTuesday, March 16, 2004 13:01:43The old Lambda +28 vdc P.S. was installed in HM on 3/11/2004.
ansanolla@naic.eduMonday, April 19, 2004 13:16:48The temp chart recorder ink pens arrived . So chart recorder back into action.
ansanolla@naic.eduThursday, May 27, 2004 12:50:51TRAK clock has been under stress due to power dips due to AEE (power company) service. It had to be adjusted a couple of times during the last two days, we have been running on emergency power for around a week.
ansanolla@naic.eduThursday, August 12, 2004 08:59:36HM is out.Is by +/- 11 milliseconds.
ansanolla@naic.eduThursday, August 12, 2004 11:48:40JPL HM was out since 8/12/2004 AST 02:01:04. DATUM clock is in since 8/12/2004 AST 10:48:00.
ansanolla@naic.eduFriday, August 13, 2004 15:57:14DATUM standard clock is shifting by 50 nsec/day vs. GPS clocks.
ansanolla@naic.eduMonday, August 16, 2004 08:43:07Calixto was setting the DATUM GPS clock outputs at back of it in order to have more 1 PPS outputs on it at 8/13/2004.
ansanolla@naic.eduThursday, August 19, 2004 15:56:10EC , JRV & ANS were working with the JPL HM in order to put it back into service. The left side vacion pump was shorted out. Still working on it.
ansanolla@naic.eduWednesday, August 25, 2004 16:07:13Bill Sisk found that the power output of the 5 mhz coming from the DATUM maser to the 5 mhz driver buffers was +17 dbm , so he add an attenuator in order to lowered to + 13 dbm.
ansanolla@naic.eduMonday, August 30, 2004 16:00:17DATUM Maser still shifting +/- 50 nsec/day.
ansanolla@naic.eduThursday, September 9, 2004 09:11:47Due to power brakedown of 1.5 hrs.(were at aux generator) clocks went bad , Bill Genter set all clocks with the DATUM GPS 4 as reference , all were less than 200 nsec.
ansanolla@naic.eduFriday, September 10, 2004 11:08:10CR & BG set the TRAK 2 clock. It was out due to power shut (1.5 hrs.) from yesterday. This clock produces the 1 PPS from DATUM maser to compare the it with the rest of the clocks.
ansanolla@naic.eduTuesday, September 14, 2004 09:15:46Personnel from Honeywell working with old HM.
ansanolla@naic.eduFriday, September 24, 2004 13:17:18Bill Genter fix the GPS 2 receiver (Joe Taylor). He found the BNC connector at the antenna output broken.
BGanolla@naic.eduWednesday, September 29, 2004 08:16:44C.Rodriguez updated firmware for NIST GPS recvr.
W. Genterwgenter@naic.eduSunday, October 31, 2004 10:38:0810/27/04 A. Nolla reported that GPS-3 was off time. On checking the software was found to be locked and not tracking. Re-initalized the software and now GPS-3 is OK.
ansanolla@naic.eduMonday, May 16, 2005 15:35:08System ok.Sometimes the program has an APT condition (Abnormal program Termination), like this month with 3 times.
ansanolla@naic.eduTuesday, July 5, 2005 09:05:23Calixto adjusted the trak 2.The Trak 1 has to be adjusted.
ansanolla@naic.eduFriday, August 5, 2005 10:54:03Bill Genter cleaned out the Joe Taylor computer disk , it was full.So it was not taking the latest data.
ansanolla@naic.eduFriday, August 26, 2005 13:13:12Clocks power was turned off by maintenance people.During a power outage at the site the electricians shut off the power braker for the timing room line conditioner.Clocks (all) were reset.
ansanolla@naic.eduMonday, August 29, 2005 10:08:33TRAK clock readind was 1.517 usec against maser. All other clocks were out , why ?
ansanolla@naic.eduTuesday, August 30, 2005 15:56:12Symmetricom maser suffer a major brake down on 8/28/2005 at AST 01:00. So the DATUM GPS 4 was set as house reference on 8/29/2005 AST 15:30.Today was found a blown fuse for the cavity heater.It will take around 2 days to stabilize according to the manufacturer advice.
ansanolla@naic.eduWednesday, August 31, 2005 11:33:41C.R. set on time the TRAK 2 receiver.Now the NIST receiver ism getting the rigth data.It was set on 8/30/2005 AST 16:07.
ansanolla@naic.eduFriday, September 2, 2005 11:22:16The Symmetricom maser back as house reference since 9/2/2005 AST 10:51. Control room notify.
ansanolla@naic.eduTuesday, September 6, 2005 14:56:29System checked after more than 72 hrs. run ,Maser stability is exellent. Still under survillance.
ansanolla@naic.eduFriday, September 30, 2005 16:14:12Yesterday Calixto replaced an UPS for our clocks.Later he reseted them.Today I found the TRAK clock out for 1 second.So I reset it & informed BG & Calixto.Clocks were ok yesterday.
ansanolla@naic.eduThursday, October 6, 2005 12:36:26On 10/4/2005n found that the antenna cable for the Don Backer GPS was bad , fix it up and GPS back to normal.Today Calixto performed a test to the TRAK 2 clock which show that the alarm for the reference input was not working.So unit is going back to factory in order to be repaired.The 1 PPS reference for sampler & NIST receiver are the outputs 9 & 10 of the 1 PPS buffers been feeded by the TRAK clock respectivily.
ansanolla@naic.eduThursday, October 6, 2005 13:54:35GPS 2 (Joe Taylor) not working.
ansanolla@naic.eduFriday, October 7, 2005 13:51:07The GPS 2 (Joe Taylor) receiver been fix replacing his burned out antenna rf section with the one of the GPS 3 , they are compatible.David Nice getting a new antenna for the GPS 2.
ansanolla@naic.eduTuesday, October 11, 2005 11:55:02TRAK 2 removed from timing room in order to check it or send it to factory for repairs.
ansanolla@naic.eduTuesday, November 29, 2005 11:24:49TRAK 2 installed today after replacing the alarm card.
ansanolla@naic.eduTuesday, January 10, 2006 13:34:47True Time clock set by Calixto.It was out by 27 usec.Set on 1/10/2006 at AST 13:15.
ansanolla@naic.eduFriday, January 13, 2006 15:37:53The antenna for the GPS 3 was hooked up.
W. Genterwgenter@naic.eduTuesday, February 7, 2006 10:17:16While conducting weekly detailed checks of the timing related equipment, the LCD display on the TrueTime Code Generator was found to be washed-out. The display intensity was at maximum. I have requested a quotation from TrueTime for a new LCD display.
ansanolla@naic.eduTuesday, March 28, 2006 09:15:48The IRIG clock at the control room was frozen ,Willy Iguina unplug it to reset it.Back to normal.
ansanolla@naic.eduMonday, May 8, 2006 10:48:33The timing room run out of power for more than 8 hours on 4/29/2006 AST 22:30 .BG came to set system.All clocks were out , Maser was ok after the 8 hours.
ansanolla@naic.eduWednesday, May 10, 2006 11:36:59The timing room power was cutoff at AST 11:08 and return it at AST 11:22 in order to replace the burned out cables of the 208 vac xformer of our power substation.Timing cabinet power (UPS)for the clocks was connected into the visitors room power.
ansanolla@naic.eduWednesday, June 21, 2006 13:13:17Manuel Alvarez replaced burned out flusforecent ligths.
ansanolla@naic.eduWednesday, July 5, 2006 10:33:21The Don Backer GPS ( #5 ) receiver removed.Going to be shipped back.
Billwgenter@naic.eduMonday, January 8, 2007 14:06:04Adjusted maser synthesizer to 57516861453 this AM. This is to reduce the heater voltage and bring the maser slope to a positive region. This will raise the frequency of the maser 1.3E-14.
ansanolla@naic.eduMonday, April 23, 2007 13:09:58On Friday 20 ,2007 we run the clocks with an external power generator feeding the UPS.The TRAK clock was reset.Rest was ok.
ansanolla@naic.eduMonday, April 23, 2007 13:11:45Forgot to mention that was due to a power shut down that was programed.
ansanolla@naic.eduThursday, May 10, 2007 13:08:53All not used rf cables were removed from timing room.Were 5 of them.
ansanolla@naic.eduWednesday, May 16, 2007 13:41:28The equipments at the timing room rack were relocated in order to make space for the clocks at the small rack on top of it on 5/14/2007.Also the batteries for the NIST receiver were replaced , it has to be done yearly.
ansanolla@naic.eduThursday, May 17, 2007 10:05:53The memory batteries of the NIST receiver were replaced yesterday , but accidentally the power cord of the receiver was removed by me , then the receiver was to be reset on its location parameters.It was done by CR , BG and RV.
Dana Whitlowdwhitlow@naic.eduThursday, November 21, 2013 16:22:28This marks the beginning of Dana's awareness that a "maser log" even existed. Some auspicious start for the official "Keeper of The Clock"!
Dana Whitlowdwhitlow@naic.eduThursday, November 21, 2013 16:27:35In late 2009 or early 2010, it was noticed that one of the the H-maser's two vacion pumps was not working. We were told that the situation could be tolerated for a long tine. Finally in late June 2011 we got around to fixing the failure. Patrick Owings from Symmetricom came out and replaced the failed pump, then the next day I replaced the other pump under Patrick's tutelage. Both replacements were accomplished without interrupting operation of the maser. As of 21 Nov 2013, both are still working normally.
Dana Whitlowdwhitlow@naic.eduThursday, November 21, 2013 16:35:21In March 2013, the IF amplitude began slowly oscillating up and down, with a period that turned out to be about 10-11 days. I contacted the Owings bros at Symmetricom, and they did not know the explanation. The Pd heater voltage was following the oscillation as if it were trying to correct or stabilize the H pressure, and it was timed nicely in phase with the If variation. There has been no discernible effect on the maser's output frequency. Now it's Nov 2013, with still no answer in sight, but still no apparent ill effect. A couple of months ago, however, Bryan Owings told me that he has since learned of another customer that experienced a similar problem, which eventually fixed itself. So we continue watchful waiting.
Dana Whitlowdwhitlow@naic.eduFriday, November 22, 2013 11:27:01NOTICE: from here on out this category should contain only items that are "pure H-maser". Other clock-related material should go in the "Clock System" category.


Back to Top